OpenStudy (anonymous):

If you flipped the graph y=x^2+2x-2 vertically, you would get the graph y=-(x^2 +2x-2) true or false

OpenStudy (anonymous):

False

OpenStudy (anonymous):

Thank-you!!

OpenStudy (anonymous):

but it was true :(

OpenStudy (shayaan_mustafa):

Hi Astemmene :) Yes it is true dear.. You are right.

OpenStudy (anonymous):

How was that true...?

OpenStudy (shayaan_mustafa):

OH wait wait.. I see mistake in my answer. In question it is whole function with -ve sign. So it should be false.

OpenStudy (shayaan_mustafa):

@Agent_Sniffles You are right. I have made mistake in seeing the question. Now I look better.

OpenStudy (anonymous):

Okay. :P I was getting confused. Haha

OpenStudy (shayaan_mustafa):

lol.. :D

OpenStudy (shayaan_mustafa):

@Astemmene Answer is false.

OpenStudy (accessdenied):

This is true. If you multiply the value of y by -1, the distance from the x-axis will not change but it will change signs. If it is positive (above the x-axis) at an x-value, then it becomes negative there. If it is negative (below the x-axis), then it will become positive. If it is 0, then it remains unchanged.

OpenStudy (shayaan_mustafa):

But whole function is multiplying with - and original is f(x)=x2+2x-2 and g(x)=-f(x) i.e. g(x)=-(x2+2x-2)

OpenStudy (shayaan_mustafa):

If it is look like this g(x)=(-x)2+2(-x)-2 then it is of the form g(x)=f(-x) and it is flipped over y-axis. i.e. vertical So it should be false. As referenced from Calculus by Howard Anton.

OpenStudy (accessdenied):

flipping over the y axis is a horizontal flip. the y-axis itself is vertical: